OpenStudy (peace420):

Driving on the highway, you can safely drive 65 miles per hour. How far can you drive in ‘h’ hours? What is the domain of the function which defines this situation? the number of hours you drive 65 the amount of gas you use the distance you drive

satellite73 (satellite73):

the domain in this case is \(h\) the hours your drive

OpenStudy (peace420):

Driving on the highway, you can safely drive 65 miles per hour. Which of the following functions could be used to find how far you can drive in ‘h’ hours? D = $ \frac {h} {65} $ D = 65 + h D = $ \frac {65} {h} $ D = 65h

satellite73 (satellite73):

those dollar signs don't work here

OpenStudy (peace420):

Okay let me try and write it a different way

satellite73 (satellite73):

no need

satellite73 (satellite73):

if you drive 65 miles per hour for two hours, how far do you go?

OpenStudy (peace420):

130 miles but thats not the question

satellite73 (satellite73):

i know it is not the question, but you answered it easily how did you get it?

OpenStudy (peace420):

i multiplied 65 by 2..

satellite73 (satellite73):

right replace the 2 with an h

OpenStudy (peace420):

D=65h??

satellite73 (satellite73):

yes as in the famous "distance equals rate times time"

OpenStudy (peace420):

Thank you Satellite!

satellite73 (satellite73):

yw

OpenStudy (peace420):

Could you help with some more questions? Please? if not thats okay